OpenStudy (firejay5):

Find the x and y intercept of the graph of each equation: 40. 2x - 6y = 12 42. 2x + 5y -10 = 0

OpenStudy (anonymous):

40. 2x-6y=12 If x is 0 then -6y = 12 y= -2 If y is 0 then 2x=12 x=6 (6,-2) 42. 2x+ 5y-10 = 0 If x is 0 then 5y=10 y=2 If y is 0 then 2x=10 x=5 (5,2)

OpenStudy (firejay5):

what's your name

OpenStudy (firejay5):

just first name